OpenStudy (anonymous):

If f(x) = x2 − 2x + 9 and g(x) = 8 − x, what is (f o g)(−4)?

OpenStudy (anonymous):

g(-4)= 8 - (-4)=12. f(g(x))= 12^2 - 2*12 +9= 129.